You are here

Beam Delivery News

Primary tabs

TItle Date Excerpt Attached images
Beam delivery update: ISAC target failure and maintenance tomorrow 01 December 2014

The cyclotron is currently off due to an ISIS trip but it was running until about 21:20, delivering ~125 µA to BL1A and 9.8 µA to BL2A. It is scheduled to deliver to BL1A and BL2A overnight, with ~4 hours maintenance to start at 08:00 tomorrow.

Beam delivery update 26 November 2014

The cyclotron’s currently delivering ~125 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue at least overnight.

Proton therapy – update 26 November 2014

There will be only four treatments after all, one each on Friday, Saturday, Sunday, and Monday. Treatments will start at ~08:00 as usual and are expected to take ~2 hours.

Beam delivery update 25 November 2014

All three beamlines are now online, and the cyclotron is currently delivering ~115 µA to BL1A, 5 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue overnight.

Beam delivery update: 4-hour maintenance tomorrow 24 November 2014

The cyclotron’s running normally, delivering ~115 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. Beam will go off at 08:30 tomorrow morning for ~4 hours of maintenance (and potentially more at ISAC to accommodate Target Hall work).

Beam delivery update: Weekend edition 21 November 2014

PIF wrapped up this afternoon and the cyclotron is delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. This is expected to continue through the weekend.

Beam delivery update 20 November 2014

The cyclotron is delivering ~110 µA to BL1A and nA as needed to BL2C1 for PIF. BL2A has been off since about 07:00 this morning for target hall work.

Beam delivery update 19 November 2014

The cyclotron is currently delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4; this is expected to continue until about 07:00, at which point BL2A will go off for ISAC target hall work.

Beam delivery update 18 November 2014

The cyclotron is currently delivering ~100 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4 for isotope production at STF; that is expected to continue overnight, with PIF resuming in the morning.

Beam delivery update 17 November 2014

The cyclotron is running and delivering ~105 µA to BL1A, 9.8 µA to BL2A, and ~95 µA to BL2C4. There is no cyclotron maintenance planned for tomorrow, but BL2A will go off around 07:00 for Target Hall work.

Pages

Subscribe to Group News